OpenStudy (anonymous):

^4 then suare root over 64x^4y^8

OpenStudy (anonymous):

is there supposed to be something before the ^4

OpenStudy (anonymous):

nope thats how it is

OpenStudy (anonymous):

at the moment this isnt really a question other than the fact that 4/64 = 1/16

OpenStudy (anonymous):

so \[\frac{ 4 }{ 64*x ^{4}*y^{8} }\] then the 4/64 can be changed to 1/16

OpenStudy (anonymous):

wait i forgot about the square root

OpenStudy (anonymous):

\[\frac{ \sqrt{4} }{ 64*x ^{4}*y ^{8} }\] \[\sqrt{4} = 2\] then 2/64 = 1/32

OpenStudy (anonymous):

that's all there is \[\frac{ 1 }{ 32*x ^{4}*y ^{8} }\]